• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name pking

    Log entry time 06:49:53 on October19,2009

    Entry number 295641

    keyword=Compton rates not stable after the detector setting changes

    The Compton rates are not stable after the changes to the detector settings (HV, table position, gain) which were made just before
    the path length adjustment.

    We do not see a clear difference betwene rates with laser on and off.

    Alex was expecting to come in shortly, and suggested that I wait for

    FIGURE 1